OpenStudy (anonymous):

a particle travels along a straight line with its acceleration at time t seconds equal to (3t+2) m/s2. the particle has an initial positive velocity and travels 30m in the fourth second. Find the velocity of the object when t=5

OpenStudy (anonymous):

\[ms^{-2}\]

OpenStudy (anonymous):

is the answer 85 m/s? Just wanna confirm before I explain because i havent done this in a while :)

OpenStudy (anonymous):

no the book says its 52 m/s :(

OpenStudy (anonymous):

hmmm im thinking lol

OpenStudy (ghazi):

see velocity = \[\int\limits_{0}^{t} a dt\]

OpenStudy (ghazi):

now integrate the function of acceleration and you'll have velocity at t=5

OpenStudy (anonymous):

haha yea i got a million answers too, but i'm going to say its 21.5 m/s. Because the acceleration is going up by 3 each second. So by the 4th second, it will be 14m/s^2. Whilst we can say the velocity at time = 4 is 7.5 m/s. So at time = 5 it should be 7.5 + 14 =21.5 m/s ..... grr